• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Adaq

    Log entry time 00:32:03 on August 9,2006

    Entry number 178468

    keyword=Shift Summary

    16:00 start with 9.5M, LD2, radiator out

    16:15 next runs might have large beam asymmetries due to tests of electron polarization measurments

    Runs:
    2361/20771 - 2365/20775 LD2, Radiator out, Polarization Measurments=>Large Beam Asymmetry

    17:05 MCC called, problems with the BSY, they cannot deliver beam to two halls simultaneously
    17:15 re-started current feed back
    17:35 beam back continue with initial program:
    17:50 lost beam, problem with a switchyard magnet

    19:40 prepare for Bulls Eye scan.
    20:00 first scan->postbone for later, switch to LH2, rad.out
    21:00 put radiator into beam
    23:00 no LN2 in Q1 right-> turn it off -> no data taking
    23:30


    Run, Events(M), Total(M), Setup
    2366/20776, 0.86, 10.3, LD2, Rad. out,
    2367/20777, 1.95, 12.25
    2368/20778, 1.95, 14.2
    2369/20779, 1.95, 16.1

    2371/207781, 0.2, 0.2 LH2 Rad out
    2372/207782, 1.95, 1.95 KH2 Rad.in
    207783
    2373/207784, 1.95, 3.9
    2374/207785, 1,95, 5.85 end of measurment